Chem 1212K-D hemistryCourse Home CHW 5 Chapter 14 Question 32-Algorithmic Constants Part A Aparticular first-order reaction has a rate oonstant of 1.35 × 10,-1 at 25.0 °C.What is the magnitude of k at 950°CifEa-555 kJ/mol. 576 O 285 x 104 O 433 x 1087 136 x 102 O 958 x 103 Submit Request Answer Provide Feedback
Answer #1

Rate constant k1=1.35x10-2 s

Temperature=T1=25oC=25+273 K=298 K

Rate constant k2=?

Temperature T2=95oC=95+273 K=368 K

Activation energy Ea=55.5 kJ/mol=55500 J/mol (1 kJ=1000 J)

We know from arrhenious equation

e RT

where k=Rate constant at temperature T

A=Frequency factor

R=Gas constant=8.314 J/Kmol

T=Temperature in K

At Temperatures T1 and T2

k_1=Ae^{\frac{-Ea}{RT_1}} and k_2=Ae^{\frac{-Ea}{RT_2}}

Dividing k2 by k1

\frac{k_2}{k_1}=\frac {e^{\frac{-Ea}{RT_2}}}{e^{\frac{-Ea}{RT_1}}}

Taking loge on both sides we get

log_e \frac{k_2}{k_1}=\frac {-Ea}{R}(\frac{1}{T_2}-\frac{1}{T_1})...equation (1)

Putting the given values in equation (1)

log_e \frac{k_2}{1.35\times10^2s^{-1}}=\frac {-55500 J}{8.314 JK^{-1}mol^{-1}}(\frac{1}{368 K}-\frac{1}{298 K})

log_e k_2 - log_e1.35\times10^2s^{-1}=\frac {-55500 Jmol^{-1}}{8.314 JK^{-1}mol^{-1}}(\frac{1}{368 K}-\frac{1}{298 K})

logek2 - 4.91=(-6675.49 K) (\frac{298-368 }{368\times298 K})

logek2 - 4.91=(-6675.49 K)(\frac{-70}{109664K})=4.26

logek2=4.26+4.91=9.17

k2=e9.17=9604.62 s-1

which is closest to the last option.

So the last option is correct.